OpenStudy (anonymous):

What is the surface area & volume of a of cylinder that has a radius of 6cm & a height of 10cm?

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cylinder}=\pi r^2h\\ \quad \\ \textit{total area of a cylinder}=2\pi r(h+r)\\ \quad \\ r= radius\qquad h=height\)

OpenStudy (nikato):

So volume of the cylinder is area of base times height. So first u would find the area of the circle

OpenStudy (polaris_s0i):

in the same respect the surface area of the cylinder is the perimeter of the circle times the height + plus the area of the 2 circles. \[2\pi rh + 2\pi r^2 = 2\pi r(r + h)\] so you're answers are \[V = \pi * 6^2 * 10 = 360\pi cm^3\] \[A = 2*\pi*6*(6 + 10) = 192\pi cm^2\]